OpenStudy (anonymous):

When Juan Pablo reaches 50 m/s a wall suddenly pops up in his lane and he hits it, coming to a complete stop in 0.35 m. What force is required to stop him at this rate?

OpenStudy (rajat97):

look the question says that juan pablo (jp) stops completely after 0.35 m after hitting the wall so when he travels 0.35 m, his velocity=0 initially , he was running with a velocity of 50m/s so we can apply the kinematical equation v^2 - u^2=2as where v=final velocity(here it is 0) u= initial velocity = (here) 50m/s s=displacement between this velocity change from u to v = (here)0.35m and a = acceleration so by putting the given values in the equation, we get a=-3571.42 m/s^2 and we know that Force=mass x acceleration so you may need the mass here if you are given the mass n the question , you candirectly multiply it with the acceleration or else i'm uploading a solution in an another way(by using impulse-momentum concept)and i think this may help hope this helps you

OpenStudy (rajat97):

and the acceleration was -ve as it is acting in the opposite direction of the velocity

OpenStudy (rajat97):

now i realise that if you do it in any way, you need the mass of the man

OpenStudy (rajat97):

@adrianf you there??